OpenStudy (anonymous):

Bernie counted the cars and motorcycles that passed his house in 45 minutes. He counted 26 vehicles and 88 tires. How many motorcycles passed his house?

OpenStudy (anonymous):

c+m=26 4c+2m=88

OpenStudy (anonymous):

use substitution method to solve

OpenStudy (anonymous):

do you know how to?

OpenStudy (anonymous):

c=26-m 4c+2m=88 Substitute c=26-m into the second equation. 4(26-m)+2m=88 Solve for m.

OpenStudy (anonymous):

the first equation c+m=26 (represents the number of cars+motorcycle that equal to 26 vehicles.) the second equation is the number of tires each vehicle has (cars have 4 tires and motorcycles have 2. Therefore, the equation came up to be 4c+2m=88) Do you kind of understand it now? :/